A survey of 539 adults aged 18-24 year olds was conducted in which they were asked what they did last Friday night. It found:
173 watched TV
176 hung out with friends
27 watched TV and ate pizza, but did not hang out with friends
41 watched TV and hung out with friends, but did not eat pizza
26 hung out with friends and ate pizza, but did not watch TV
39 watched TV, hung out with friends, and ate pizza
73 did not do any of these three activities
How may 18-24 year olds (of these three activities) only ate pizza last Friday night?

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13209)



(1) 173 watched TV
(2) 176 hung out with friends
(3) 27 watched TV and ate pizza, but did not hang out with friends
(4) 41 watched TV and hung out with friends, but did not eat pizza
(5) 26 hung out with friends and ate pizza, but did not watch TV
(6) 39 watched TV, hung out with friends, and ate pizza
(7) 73 did not do any of these three activities

Draw a Venn diagram of three mutually intersecting circles, one for each activity.

Fill in the information (3) to (6).

Use diagram along with (1) and (2) to find the numbers that only watched TV or that only hung out with friends.

Use all the numbers in the diagram, along with (7) and the total number of 539, to find the number that only ate pizza.


Answer by ikleyn(52878)

.
A survey of 539 adults aged 18-24 year olds was conducted in which they were asked what they did last Friday night. It found:
173 watched TV
176 hung out with friends
27 watched TV and ate pizza, but did not hang out with friends
41 watched TV and hung out with friends, but did not eat pizza
26 hung out with friends and ate pizza, but did not watch TV
39 watched TV, hung out with friends, and ate pizza
73 did not do any of these three activities
How may 18-24 year olds (of these three activities) only ate pizza last Friday night?
~~~~~~~~~~~~~

We are given a universal set U of 539 persons and its subsets

T  (watched TV)                                                          of 173

F  (hung out with friends)                                               of 176

TP \ TPF  (watched TV and ate pizza, but did not hang out with friends)  of  27

TF \ TPF  (watched TV and hung out with friends, but did not eat pizza)  of  41

FP \ TPF  (hung out with friends and ate pizza,  but did not watch TV)   of  26 

TPF       (watched TV, hung out with friends, and ate pizza)             of  39

The complement of (T U F U P) to U                                       of  73.


Couple comments regarding this table.

    For brevity, I write AB for the intersection of subsets A and B (instead of traditional (A ∩ B) ).
    For brevity, I write TPF for the triple intersection (T ∩ P ∩ F).
    The sign  " \ "  means subtraction of subsets.



Now, in the elementary set theory, there is a REMARCABLE identity 

    n(A U B U C) = n(A) + n(B) + n(C) - n(AB) - n(AC) - nBC) + n(ABC)              (1)

for any three subsets A, B, C of a universal set, their in-pair intersections AB, AC and BC, and the triple intersection ABC.



I will apply it for the given subsets T, F, P, their in-pairs and triple intersetions

    n(T U F U P) = n(T) + n(F) + n(P) - n(TF)   - n(TP)   - n(FP)   + n(TPF).      (2)



In this equality, all the terms are given and are known, EXCEPT only one n(P).  It is my goal now to find n(P).
For it, I substitute all the known values into equation (2)

    539 - 73     = 173  + 176  + n(P) - (41+39) - (27+39) - (26+39) + 39.          (3)



From (3) I get

    n(P) = (539-73) - 173 - 176 + (41+39) + (27+39) + (26+39) - 39 = one click in my Excel = 289.


Now my last step is

    the number of those who only eat Pizza = n(P) - n(TP) - n(FP) + n(TPF) = 289 - (27+39) - (26+39) + 39 = 197.


ANSWER.  The number of those who only eat Pizza is 197.

So, the problem is just solved, and the answer is 197 (by the modulus of my possible arithmetic errors).


' ' ' ' ' ' ' '

I like to get  "THANKS"  from students for my work and for my teaching.

So,  I expect to get your  "DEEPEST THANKS"  soon.